OpenStudy (usukidoll):

Can anyone give me a step by step tutorial on how to solve an equation with an undetermined coefficient? Ex. (D^3-D)y=2e^x I know that I can factor out a D (D)(D^2-1)y=2e^x and I can factor again (D)(D+1)(D-1)y=2e^x but I don't know what to do next.

OpenStudy (usukidoll):

@Euler271

OpenStudy (usukidoll):

well there is an exponential on the right... and that's just one.. so replace the one from e^x in the D?!??!?!

OpenStudy (anonymous):

i have no idea what the (D - x) notation is. the way i'd do it is: let y = C_1e^x + C_2xe^x so (D^3 - D)y = 2e^x becomes \[[(C_1e^x + C_2(x+3)e^x) - C_1e_x - C_2(x+1)e^x] = 2e^x\]\[2C_2e^x = 2e^x\] so C_1 = 0 and C_2 = 1 making the particular solution: y = xe^x hope this helps. wish i knew the notation lol

zepdrix (zepdrix):

I think it's just a nice simple notation for the derivative operator :)\[\Large (D^3-D)y\quad=\quad y'''-y'\]

OpenStudy (usukidoll):

it's another format but it's similar to y'''+y'''+y''+y'+y=0

OpenStudy (anonymous):

ya. i know that about it, but not looking at it in its factored form and getting info from it

OpenStudy (anonymous):

im sure there is a simpler way to get solutions of the des just by looking at the factored form

OpenStudy (usukidoll):

if the right hand side was = 0, we would've been done, but that's not the case here

OpenStudy (anonymous):

that is right. the total solution is the homogeneous solution [when right side = 0] + particular solution [when right side =/= 0]. sometimes the total solution is only the homogeneous part. only when = 0

OpenStudy (anonymous):

to get the rest of the solution, you'd pretend the right side = 0

OpenStudy (usukidoll):

why am I letting y be let y = C_1e^x + C_2xe^x??

OpenStudy (usukidoll):

o...pretend that it's 0 well that would be let's see D=0,1,-1 c1+c2e^x+c3e^-x=y

OpenStudy (anonymous):

i think i get the notation now lol. but for my choice of y: that is the way i'd do it and that would be a method of undetermined coefficients. if the differential equation were D^3 - D = e^(2x), i would have chosen the form y = Ce^(2x) and played it out, giving: 8Ce^(2x) - 2e^(2x) = e^(2x) ----> C = 6 ----> y_particular = 6e^(2x) But in our case, the right side [the particular part, aka the forcing function], taking just Ce^x leads to Ce^x - Ce^x = 2e^x = 0 so we can't solve for C. that is why i applied the common trick of using a close form which will lead in being able to determine the coefficient(s). so i chose, by habit, C_1e^x + C_2xe^x

OpenStudy (usukidoll):

oh gawd I'm confused x(

OpenStudy (usukidoll):

so.... ok.. if maybe D = 2e^x then